# MSCN: Multi-view Structural Convolution Network for Domain-Invariant Point Cloud Recognition of Autonomous Vehicles

### v4

@@ Abstract
Although LiDAR sensors have become indispensable for autonomous vehicles (AVs) due to their ability to provide accurate 3D scene understanding and robust perception under adverse weather conditions, the properties of LiDAR point clouds vary widely across sensor configurations and data acquisition domains, leading to severe performance degradation when models are transferred between heterogeneous sensors or from simulation to the real world. To address this challenge, we propose the Multi-view Structural Convolution Network (MSCN), a novel architecture designed to achieve domain-invariant recognition across diverse LiDAR configurations and environments. MSCN comprises Structural Convolution Layers (SCL) that extract local context geometric features from point clouds and Structural Aggregation Layers (SAL) that extract and aggregate both local and overall context features from point clouds. Furthermore, we incorporate an unseen domain generation strategy to mitigate domain gaps during training. Extensive experiments demonstrate that MSCN consistently outperforms state-of-the-art point cloud classification methods across all domain change scenarios. These results highlight MSCN as a scalable solution for deploying LiDAR-based perception systems of AVs. Our code is available at https://github.com/MLMLab/MSCN.
Index Terms:
Point cloud-based Autonomous Vehicle System, LiDAR Point Cloud Classification, Domain-Invariant Feature Representation, Unseen Domain Generation, Domain Generalization.
@@ I Introduction
Autonomous vehicles (AVs) are envisioned as a transformative solution to major transportation challenges, including the mitigation of traffic accidents [1, 2], congestion alleviation [3, 4], and improved efficiency in transportation systems [5]. Reliable perception is central to enabling these benefits, as AVs must accurately recognize and interpret their surroundings to ensure safety and efficiency. In this context, Light Detection and Ranging (LiDAR) has emerged as a particularly valuable sensor because it provides rich geometrical information, such as depth and structural details, that complements traditional camera-based perception. LiDAR-based point clouds provide accurate geometric representations of traffic environments, offering precise depth and structural information that is indispensable for robust AV perception and intelligent transportation systems (ITS).
@@
Despite the advances of LiDAR-based studies, prior research has largely overlooked the importance of achieving robust perception across heterogeneous LiDAR sensors. In practice, as illustrated in Figure 1, the properties of point clouds vary significantly depending on both the sensor configuration and the data acquisition domain. First, different sensor configurations, such as variations in the number of LiDAR channels, cause substantial discrepancies in point cloud density and resolution. When models are trained on high-resolution LiDARs but deployed on cost-efficient low-resolution LiDARs, they often experience severe performance degradation. Since LiDAR configurations vary widely across manufacturers and even across research platforms, this mismatch presents a critical barrier to the scalable adoption of autonomous driving systems. Second, point cloud characteristics vary across geographic locations, as object appearances and structural patterns differ in distinct environments. Without accounting for such location-specific variations, autonomous vehicles risk misinterpreting unfamiliar object shapes, which can lead to failures in recognition and ultimately compromise driving safety. Moreover, the widely recognized sim-to-real gap poses another major challenge: while simulation environments enable the generation of abundant data and diverse accident scenarios, the point clouds they produce are typically uniform and noise-free. By contrast, real-world LiDAR data are often sparse and irregular, resulting in a pronounced distributional discrepancy that limits the generalization of models trained solely on simulation data.
@@
The main contributions of this study can be summarized as follows:
@@
  • We design progressively harder experimental settings to rigorously evaluate the domain-invariant recognition properties of MSCN: (1) uniform synthetic point clouds with controlled geometric perturbations, (2) indoor real-world point clouds with geometric variations, (3) autonomous vehicle datasets with varying LiDAR channel configurations and geographic locations, (4) sim-to-real domain shifts from synthetic to real driving datasets, and (5) combined challenges of channel change, geographic variation, and geometric perturbations in autonomous vehicle datasets. These progressively harder settings demonstrate MSCN’s robust geometric-invariant recognition properties, which are essential for ensuring reliable perception in AVs under diverse deployment conditions.
@@ II-A1 Multi-view-based methods
Considering that a collection of 2D views can offer significant information for 3D shape recognition compared to traditional manual feature extraction, several studies have explored recognizing 3D shapes from collections of rendered 2D images [26], [27], [28], [29]. Inspired by these works, Hamdi et al. proposed the Multi-View Transition Network (MVTN) [30]. This network includes a differentiable module that predicts the optimal viewpoint for a task-specific multi-view network, aiming to enhance 3D shape recognition by addressing limitations of previous multi-view methods. However, multi-view approaches still struggle to fully exploit the geometric relationships inherent in 3D data and extend their applicability to other 3D tasks, such as segmentation and reconstruction.
@@ II-A2 Voxel-based methods
To manage the unstructured nature of point clouds, some studies have used voxel grids or voxelized 3D shapes as model inputs, enabling the application of standard 3D CNNs [31], [32], [33]. Researchers like Riegler et al. [34] and Wang et al. [35] have employed octree structures instead of voxels to improve computational efficiency and reduce memory consumption. Despite their focus on the relationships within 3D data and the grouping of point clouds, voxel-based methods often suffer from low classification efficiency due to the inherent sparsity and incomplete information in point clouds.
@@ II-C1 Domain adaptation
Recent studies have demonstrated that deep learning models are prone to significant performance degradation when evaluated on out-of-distribution (OOD) datasets, even with minor variations in the data generation process [17], [18]. These studies underscore the limited generalization capacity of deep learning models, which can severely impact their performance in real-world applications. Domain adaptation (DA) has emerged as a strategy to tackle the OOD data challenge by utilizing a subset of data from the target domain [19], [20], [21]. This approach enables a model trained on a source domain with different distributions to adapt to the target domain. The fundamental concept of DA is that, despite differences between source and target domains, common underlying structures or patterns can be leveraged to enhance model performance in the target domain. However, DA assumes the availability of either labeled or unlabeled data from the target domain for model adaptation, a condition not always met in real-world scenarios. Acquiring labeled data from the target domain can be prohibitively time-consuming, expensive, or impractical due to ethical or legal constraints. Additionally, the target domain’s data distribution may remain unclear, complicating the application of conventional DA methods that rely on understanding this distribution.
@@ II-C2 Single domain generalization
To address the challenges associated with DA, many domain generalization (DG) studies, particularly those focusing on single domain generalization, have been proposed [22], [23], [24], [25]. Single domain generalization is a particularly challenging form of model generalization where models are trained on a single domain and tested on unseen domains. A promising approach involves learning cross-domain invariant representations by expanding the coverage of the training domain. These methods typically synthesize additional samples in feature space to broaden the data distribution in the training set. Zhang et al. [22] observed that differences in medical images mainly arise in three aspects: image quality, image appearance, and spatial configuration. They augment these aspects through data augmentation. However, this method requires selecting the appropriate type and magnitude of augmentation based on the target domain, which can be challenging for other tasks. Volpi et al. [23] and Qiao et al. [24] synthesize additional data through adversarial learning to enhance model robustness. Nevertheless, this approach has limitations: the types of augmentation are relatively simple, and using too many adversarial examples during training can degrade classifier performance. To mitigate these issues, Li et al. [25] proposed progressively generating multiple domains and learning a feature extractor using both source and generated data. The model employs contrastive learning to develop domain-invariant representations, effectively clustering each class to facilitate better decision boundaries and improve generalization. While this approach equips models to perform well in real-world scenarios, particularly on previously unseen domains, it is currently only applicable to domain generalization tasks involving well-structured 2D images.
III Multi-view Structural Convolution Network (MSCN)
@@ III-A Structural Convolution Layer (SCL)
Figure 2: Illustration of the Multi-view Structural Convolution Network (MSCN) architecture, including (a) the classification architecture, (b) the Structural Convolution Layer, and (c) the Structural Aggregation Layer.
@@
Using the defined kernels, we perform convolution on irregular points within a receptive field of a point cloud. Specifically, we first assess the similarity between the features within the receptive field of $p_{n}$ (i.e. $F(p_{n})$, $F(p_{m})$ $\forall$ $p_{m}\in N(p_{n},M)$, as specified in Equation 1) and the weight vectors of the directional kernel $K^{S}_{dir}$, centered around $b^{C}_{dir}$ with $S$ branches (namely, $w(b^{c}_{dir})$, $w(b^{s}_{dir})$ $\forall$ $s=1,2,\ldots,S$). We consider every pairing of $(p_{m},b^{s}_{dir})$. Consequently, the direction-based convolution between a receptive field and a direction-based kernel can be described as:
@@ III-B Structural Aggregation Layer (SAL)
where $P_{in}\in\mathbb{R}^{N_{i-1}\times 3}$$P_{out}\in\mathbb{R}^{N_{i}\times 3}$, $F_{in}\in\mathbb{R}^{N_{i-1}\times D_{i-1}}$, $F_{out}\in\mathbb{R}^{N_{i}\times D_{i-1}}$, and $N_{i}=\frac{N_{i-1}}{r}$. This local feature aggregation operation allows us to learn multi-scale local context features, which are crucial factors for extracting local geometric information.
@@ III-C Classification Process
IV PROGRESSIVE DOMAIN EXPANSION PROCESS
@@ IV-A Generator Composition
Figure 3: Progressive domain expansion process, including (a) the modified MSCN for contrastive learning and (b) the training architecture with the progressive domain expansion process.
@@ V EVALUATION
To evaluate the domain-invariant capabilities of the proposed MSCN, we conducted comprehensive experiments using both synthetic datasets and real-world datasets collected from autonomous vehicles. The synthetic dataset experiments aim to verify MSCN’s invariance to geometric transformations, including translation and scaling. These tests provide insights into the model’s ability to maintain performance under controlled yet challenging conditions.
In the real-world dataset experiments, we evaluated MSCN’s robustness in scenarios where the LiDAR resolution differs between training and testing environments. This setup simulates practical conditions in autonomous driving, where sensor configuration may vary. Furthermore, we assessed MSCN’s performance under severe domain shifts, combining both dataset-specific variations and additional transformations such as rotation and translation. Even in these harsh conditions, our MSCN consistently demonstrates robust performance, highlighting its effectiveness in domain-adaptive tasks.
V-A Dataset and Implementation Details
@@ V-A1 Synthetic Datasets
We conduct experiments using ShapeNetPart [48] and ModelNet40 [38] datasets. For the ShapeNetPart dataset, we evaluate the proposed MSCN on the part segmentation task, while for the ModelNet40 dataset, we perform object classification. Both experiments followed the standard training settings commonly used in previous work. In addition, specific experiments are conducted comparing MSCN with PointMLP [15] and PointTransformer [47] to demonstrate MSCN’s invariance to geometric transformations such as translation and scaling. For these experiments, we fixed the number of training epochs for all models to 30 and used a consistent point cloud size of 1024 points. These settings ensure a fair comparison across models and highlight MSCN’s robustness under varying geometric transformations.
Figure 4: Qualitative examples of geometric transformation on ShapeNetPart[32], where scale 10 represents a transformation where the original point cloud is scaled by a factor of 10, while shift 90 indicates a random translation of the point cloud by a distance of 90 units.
TABLE I: Object part segmentation results on ShapeNetPart[32]. mIOU values are reported for each method.
@@
MSCN (Ours) 1024 84.8
V-A2 Real-world Dataset
V-A3 Real-World Autonomous Vehicle Datasets
We conduct experiments using three real-world point cloud datasets, namely KITTI[36], PanKyo[41], and nuScenes[37], as well as a synthetic dataset from the MORAI simulator[42]. These datasets, accessible through the following link  [^1] https://sites.google.com/site/cvsmlee/dataset, originate from various countries and manufacturers, making them crucial for evaluating the inference capacity of the model and the robustness to domain shifts.
Comparisons between our methods are conducted, including MSCN and MSCN, and precious methods including PointNet[12], PointNeXt[40], PointMLP[15], DGCNN[14], GCN[16], SCN[41], PointTransformer[47], and PointMamba[46]. PointMamba was first pre-trained on ShapeNet[48] and then trained on KITTI, PanKyo, and nuScenes for the classification task. The networks are evaluated through the a standard accuracy metric.
To prevent overfitting, which could lead to incorrect classification performance in domain change scenarios, the number of training epochs for all methods, except MSCN, is limited to 10. MSCN indicates that applying the unseen domain generation methods to MSCN. For MSCN, a new domain is generated every 15 training cycles and this process is repeated 20 times, resulting in a total of 300 epochs.
We first conduct experiments on the ShapeNetPart [32] dataset for part segmentation. Table I summarizes the part segmentation results, where mIoU values are reported for each method. Our MSCN achieves performance comparable to state-of-the-art (SOTA) models, despite using only 1024 points, unlike other SOTA models that rely on 2048 points. While SOTA models exhibit high performance on the given dataset, MSCN’s performance is slightly inferior compared to SOTA models. This performance gap can be explained by the inherent trade-off between robustness and accuracy, as discussed in [68]. They showed that models performing robustly often lead to a shift in focus from minimizing standard classification errors (natural error) to improving robustness, which can degrade accuracy in the ideal condition (e.g. same properties between training and test dataset). Note that our MSCN were designed to generalize well to harsh or out-of-distribution scenarios while preserving accuracy on natural data. To evaluate the robustness of MSCN, we conduct experiments to visualize the robustness of MSCN compared to PointMLP and PointTransformer under geometric transformations. Figure 4 illustrates the results of scaling and translating point clouds. MSCN demonstrates robust part segmentation even under significant geometrical transformations, whereas PointMLP and PointTransformer show sensitivity to these changes, resulting in suboptimal performance. These results emphasize MSCN’s ability to generalize effectively across various geometric variations.
Figure 5: Evaluation of geometrical invariance properties on ModelNet40 [38], showing model accuracies under translation (0–200 units) and scaling (0.01–100).
@@ V-B Evaluation on Synthetic Datasets: Robustness to Geometric Deformation
Furthermore, we conduct experiments on the ModelNet40 [38] dataset for object classification. Table II presents the classification results, where MSCN is slightly lower than some SOTA models. Although SOTA models demonstrate strong performance under ideal conditions, their sensitivity to geometric transformations can result in significant performance degradation in more challenging scenarios. To further investigate this, Figure 5 illustrates the performance of models under geometric transformations such as translation and scaling. MSCN consistently exhibits robust performance across all transformations, maintaining stable accuracy despite severe geometric changes. In contrast, other methods experience substantial drops in accuracy, highlighting their vulnerability to such transformations. These results underscore the strength of MSCN in terms of geometrical invariance, demonstrating its potential for real-world applications where robustness to geometric variations is critical.
However, it is important to note that ModelNet40 and ShapeNetPart are derived from Computer Aided Design (CAD), which are generated in highly idealized environments. In contrast, real-world point clouds are often sparse and only capture surface-level details, which makes them fundamentally different from synthetic datasets. To address this disparity, we extend our experiments to real-world point clouds, enabling a more comprehensive evaluation of our MSCN’s performance in practical scenarios.
TABLE III: Classification results on the ScanObjectNN[67] dataset. PT indicates Pretrained models, RT represents Random Translation, and RS denotes Random Scaling.
@@
TABLE IV: Intra-domain classification results on AV datasets.

### v5

# Multi-view Structural Convolution Network for Domain-Invariant Point Cloud Recognition of Autonomous Vehicles

@@ Abstract
Point cloud representation has recently become a research hotspot in the field of computer vision and has been utilized for autonomous vehicles. However, adapting deep learning networks for point cloud data recognition is challenging due to the variability in datasets and sensor technologies. This variability underscores the necessity for adaptive techniques to maintain accuracy under different conditions. In this paper, we present the Multi-View Structural Convolution Network (MSCN) designed for domain-invariant point cloud recognition. MSCN comprises Structural Convolution Layers (SCL) that extract local context geometric features from point clouds and Structural Aggregation Layers (SAL) that extract and aggregate both local and overall context features from point clouds. Furthermore, MSCN enhances feature robustness by training with unseen domain point clouds generated from the source domain, enabling the model to acquire domain-invariant representations. Extensive cross-domain experiments demonstrate that MSCN achieves an average accuracy of 82.0%, surpassing the strong baseline PointTransformer by 15.8%, confirming its effectiveness under real-world domain shifts. Our code is available at https://github.com/MLMLab/MSCN.
@@ I Introduction
Autonomous vehicles (AVs) are envisioned as a transformative solution to major transportation challenges, including the mitigation of traffic accidents [1, 2, 3], congestion alleviation [4, 5], and improved efficiency in transportation systems [6]. Reliable perception is central to enabling these benefits, as AVs must accurately recognize and interpret their surroundings to ensure safety and efficiency. In this context, Light Detection and Ranging (LiDAR) has emerged as a particularly valuable sensor because it provides rich geometrical information, such as depth and structural details, that complements traditional camera-based perception. LiDAR-based point clouds provide accurate geometric representations of traffic environments, offering precise depth and structural information that is indispensable for robust AV perception and intelligent transportation systems (ITS).
@@
Despite the advances of LiDAR-based studies, prior research has largely overlooked the importance of achieving robust perception across heterogeneous LiDAR sensors. In practice, as illustrated in Figure 1, the properties of point clouds vary significantly depending on both the sensor configuration and the data acquisition domain. First, different sensor configurations, such as variations in the number of LiDAR channels, cause substantial discrepancies in point cloud density and resolution. For example, when models are trained on high-resolution LiDARs but deployed on cost-efficient low-resolution LiDARs, they often experience severe performance degradation. Since LiDAR configurations vary widely across manufacturers and even across research platforms, this mismatch presents a critical barrier to the scalable adoption of autonomous driving systems. Second, point cloud characteristics vary across geographic locations, as object appearances and structural patterns differ in distinct environments. Without accounting for such location-specific variations, autonomous vehicles risk misinterpreting unfamiliar object shapes, which can lead to failures in recognition and ultimately compromise driving safety. Moreover, the widely recognized sim-to-real gap poses another major challenge: while simulation environments enable the generation of abundant data and diverse accident scenarios, the point clouds they produce are typically uniform and noise-free. In contrast, real-world LiDAR data are often sparse and irregular, resulting in a pronounced distributional discrepancy that limits the generalization of models trained solely on simulation data.
@@
The main contributions of this study can be summarized in three folds as follows:
@@
  • We design progressively harder experimental settings to rigorously evaluate the domain-invariant recognition properties of MSCN: (1) uniform synthetic point clouds with controlled geometric perturbations, (2) indoor real-world point clouds with geometric variations, (3) autonomous vehicle datasets with varying LiDAR channel configurations and geographic locations, (4) sim-to-real domain shifts, and (5) combined challenges of channel change and geometric perturbations. Through (1)–(2), we reveal that existing models remain highly sensitive to geometric transformations even under idealized synthetic and indoor environments. Evaluations on autonomous driving datasets (3)–(5) further demonstrate MSCN’s robustness under both real-to-real and sim-to-real domain shifts, achieving an average cross-domain accuracy of 82.0%, surpassing the strong baseline PointTransformer by 15.8%.
@@ II-A1 Multi-view-based methods
Considering that a collection of 2D views can offer significant information for 3D shape recognition compared to traditional manual feature extraction, several studies [27], [28], [29], [30] have explored recognizing 3D shapes from collections of rendered 2D images. Inspired by these works, Hamdi et al. [31] proposed the Multi-View Transition Network (MVTN). This network includes a differentiable module that predicts the optimal viewpoint for a task-specific multi-view network, aiming to enhance 3D shape recognition by addressing limitations of previous multi-view methods. However, multi-view approaches still struggle to fully exploit the geometric relationships inherent in 3D data and extend their applicability to other 3D tasks, such as segmentation and reconstruction.
@@ II-A2 Voxel-based methods
To manage the unstructured nature of point clouds, some studies [32], [33], [34] have used voxel grids or voxelized 3D shapes as model inputs, enabling the application of standard 3D CNNs. Researchers like Riegler et al. [35] and Wang et al. [36] have employed octree structures instead of voxels to improve computational efficiency and reduce memory consumption. Despite their focus on the relationships within 3D data and the grouping of point clouds, voxel-based methods often suffer from low classification efficiency due to the inherent sparsity and incomplete information in point clouds.
@@ II-B1 Point-by-point methods
Beyond classification, several recent works have focused on robust feature representation for registration tasks, which are closely related to our goal of learning domain-invariant geometric features. Ao et al. [16] introduced BUFFER, a point-wise learner designed to predict keypoints and estimate point orientations, achieving a balance among accuracy, efficiency, and generalizability in registration. She et al. [17] proposed PosDiffNet, which leverages a graph-neural PDE based on Beltrami flow and integrates position embeddings into a Transformer-based ODE framework for multi-level correspondence learning, achieving state-of-the-art performance under large fields of view with perturbations.
@@ II-C1 Domain adaptation
Recent studies [18], [19] have demonstrated that deep learning models are prone to significant performance degradation when evaluated on out-of-distribution (OOD) datasets, even with minor variations in the data generation process. These studies underscore the limited generalization capacity of deep learning models, which can severely impact their performance in real-world applications. Domain adaptation (DA) [20], [21], [22] has emerged as a strategy to tackle the OOD data challenge by utilizing a subset of data from the target domain. This approach enables a model trained on a source domain with different distributions to adapt to the target domain. The fundamental concept of DA is that, despite differences between source and target domains, common underlying structures or patterns can be leveraged to enhance model performance in the target domain. However, DA assumes the availability of either labeled or unlabeled data from the target domain for model adaptation, a condition not always met in real-world scenarios. Acquiring labeled data from the target domain can be prohibitively time-consuming, expensive, or impractical due to ethical or legal constraints. Additionally, the target domain’s data distribution may remain unclear, complicating the application of conventional DA methods that rely on understanding this distribution.
[…]
Figure 2: The Multi-view Structural Convolution Network (MSCN) architecture for 3D point cloud classification comprises several components. (a) The MSCN includes feature extraction layers, specifically Structural Convolution Layers (SCL) and Structural Aggregation Layers (SAL), as well as Global Max Pooling and a multi-layer perceptron (MLP) for classification. (b) SCL is engineered to extract local features from each point. (c) The SAL is designed to combine local context features with overall context features.
@@ II-C2 Single domain generalization
To address the challenges associated with domain adaptation (DA), many domain generalization (DG) studies [23], [24], [25], [26] particularly those focusing on single domain generalization, have been proposed. Single domain generalization is a particularly challenging form of model generalization where models are trained on a single domain and tested on unseen domains. A promising approach involves learning cross-domain invariant representations by expanding the coverage of the training domain. These methods typically synthesize additional samples in feature space to broaden the data distribution in the training set. Zhang et al. [23] observed that differences in medical images mainly arise in three aspects: image quality, image appearance, and spatial configuration. They augment these aspects through data augmentation. However, this method requires selecting the appropriate type and magnitude of augmentation based on the target domain, which can be challenging for other tasks. Volpi et al. [24] and Qiao et al. [25] synthesize additional data through adversarial learning to enhance model robustness. Nevertheless, this approach has limitations: the types of augmentation are relatively simple, and using too many adversarial examples during training can degrade classifier performance. To mitigate these issues, Li et al. [26] proposed progressively generating multiple domains and learning a feature extractor using both source and generated data. The model employs contrastive learning to develop domain-invariant representations, effectively clustering each class to facilitate better decision boundaries and improve generalization. While this approach equips models to perform well in real-world scenarios, particularly on previously unseen domains, it is currently only applicable to domain generalization tasks involving well-structured 2D images.
III Multi-view Strucutral Convolution Network (MSCN)
@@ III-A Structural Convolution Layer (SCL)
Using the defined kernels, we perform convolution on irregular points within a receptive field of a point cloud. Specifically, we first assess the similarity between the features within the receptive field of $p_{n}$ (i.e. $F(p_{n})$, $F(p_{m})$ $\forall$ $p_{m}\in N(p_{n},M)$, as specified in Equation 1 and the weight vectors of the directional kernel $K^{S}_{dir}$, centered around $b^{C}_{dir}$ with $S$ branches (namely, $w(b^{c}_{dir})$, $w(b^{s}_{dir})$ $\forall$ $s=1,2,\ldots,S$). We consider every pairing of $(p_{m},b^{s}_{dir})$. Consequently, the direction-based convolution between a receptive field and a direction-based kernel can be described as:
@@ III-B Structural Aggregation Layer (SAL)
where $P_{in}\in\mathbb{R}^{N_{i-1}\times 3}$, $P_{out}\in\mathbb{R}^{N_{i}\times 3}$, $F_{in}\in\mathbb{R}^{N_{i-1}\times D_{i-1}}$, $F_{out}\in\mathbb{R}^{N_{i}\times D_{i-1}}$, and $N_{i}=\frac{N_{i-1}}{r}$. This local feature aggregation operation allows us to learn multi-scale local context features, which are crucial factors for extracting local geometric information.
[…]
Figure 3: The architecture for generating unseen domains in MSCN to learn more robust feature representations includes the following components. In (a), the modified MSCN is shown, consisting of a feature extractor (F), a classification head (C), and a projection head (P) with the projection head added for contrastive learning. In (b), the process of generating arbitrary unseen domains and training the MSCN with both source and generated data is illustrated. These generation and training processes are performed alternately.
@@ III-C Classification Process
IV Progressive Domain Expansion Progress
@@ V Evaluation
To evaluate the domain-invariant capabilities of the proposed MSCN, we conducted comprehensive experiments using both synthetic datasets and real-world datasets collected from self-driving vehicles. The synthetic dataset experiments aim to verify MSCN’s invariance to geometric transformations, including translation and scaling. These tests provide insights into the model’s ability to maintain performance under controlled yet challenging conditions.
In the real-world dataset experiments, we evaluated MSCN’s robustness in scenarios where the LiDAR resolution differs between training and testing environments. This setup simulates practical conditions in autonomous driving, where sensor configurations may vary. Furthermore, we assessed MSCN’s performance under severe domain shifts, combining both dataset-specific variations and additional transformations such as rotation and translation. Even in these harsh conditions, our MSCN consistently demonstrates robust performance, highlighting its effectiveness in domain-adaptive tasks.
V-A Datasets & Implementation Detail
@@ V-A1 Synthetic Datasets
We conducted experiments using ShapeNetPart [49] and ModelNet40 [39] datasets. For the ShapeNetPart dataset, we evaluated the proposed MSCN on the part segmentation task, while for the ModelNet40 dataset, we performed object classification. Both experiments followed the standard training settings commonly used in previous work. In addition, specific experiments are conducted comparing MSCN with PointMLP [14] and PointTransformer [48] to demonstrate MSCN’s invariance to geometric transformations such as translation and scaling. For these experiments, we fixed the number of training epochs for all models to 30 and used a consistent point cloud size of 1024 points. These settings ensured a fair comparison across models and highlighted MSCN’s robustness under varying geometric transformations.
V-A2 Real-world Indoor Dataset
V-A3 Real-World Self-Driving Vehicle Datasets
We conducted experiments using three real-world point cloud datasets, namely KITTI [37], PanKyo [42], and nuScenes [38], as well as a synthetic dataset from the MORAI simulator [43]. These datasets, accessible through the following link  [^1] https://sites.google.com/site/cvsmlee/dataset, originate from various countries and manufacturers, making them crucial for evaluating the inference capacity of the model and the robustness to domain shifts.
Comparisons between our methods are conducted, including MSCN and MSCN, and precious methods including PointNet [11], PointNeXt [41], PointMLP [14], DGCNN [13], GCN [15], SCN [42], PointTransformer [48], and PointMamba [47]. PointMamba was first pre-trained on ShapeNet [49] and then trained on KITTI, PanKyo, and nuScenes for the classification task. The networks were evaluated through the a standard accuracy metric.
To prevent overfitting, the number of training epochs for all methods, except MSCN, is limited to 10. MSCN indicates that applying the unseen domain generation methods to MSCN. For MSCN, a new domain is generated every 15 training cycles and this process is repeated 20 times, resulting in a total of 300 epochs.
TABLE I: Object part segmentation results on ShapeNetPart [33], grouped by dataset. mIOU values are reported for each method and dataset.
@@ V-B Evaluation on Synthetic Datasets: Robustness to geometric deformation
MSCN(Ours) 2048 85.0
[…]
Figure 4: Examples of geometric transformation on ShapeNetPart [33], where scale 10 represents a transformation where the original point cloud is scaled by a factor of 10, while shift 90 indicates a random translation of the point cloud by a distance of 90 units.
We first conducted experiments on the ShapeNetPart [33] dataset for part segmentation. Table I summarizes the part segmentation results, where mIoU values are reported for each method. While SOTA models exhibit high performance on the given dataset, MSCN’s performance is slightly inferior compared to SOTA models. This performance gap can be explained by the inherent trade-off between robustness and accuracy, as discussed in [70]. Our MSCN were designed to generalize well to harsh or out-of-distribution scenarios while preserving accuracy on natural data. However, Zhang et.al [70] showed this approach often leads to a shift in focus from minimizing standard classification errors (natural error) to improving robustness, which can degrade accuracy in ideal.
[…]
[width=0.95]figures/ModelNet40_invariance_property.png
Figure 5: Evaluation of Geometrical Invariance Properties on ModelNet40 [39]. Each graph illustrates the accuracy of various models under specific geometric transformations applied to point clouds. (a) Translation: Accuracy as the translation magnitude increases sequentially from 0 to 200 units. (b) Scaling: Accuracy as the scale factor increases sequentially from 0.01 to 100.
To further evaluate robustness of MSCN, we conducted experiments to visualize the robustness of MSCN compared to PointMLP and PointTransformer under geometric transformations. Figure 4 illustrates the results of scaling and translating point clouds. MSCN demonstrates robust part segmentation even under significant geometrical transformations, whereas PointMLP and PointTransformer show sensitivity to these changes, resulting in suboptimal performance. These results emphasize MSCN’s ability to generalize effectively across various geometric variations.
@@
TABLE III: Classification results on the ScanObjectNN [69] dataset. This table summarizes the performance of methods under different experimental settings, including self-supervised pretraining, supervised learning, and the evaluation of MSCN’s translation and scale invariance. PT, RT, and RS indicates Pretrained models, Random Translation, and Random Scaling, respectively.
@@
Furthermore, we conducted experiments on the ModelNet40 [39] dataset for object classification. Table II presents the classification results, where MSCN is slightly lower than some SOTA models. Although SOTA models demonstrate strong performance on this dataset, their sensitivity to geometric transformations can result in significant performance degradation in more challenging scenarios. To further investigate this, Figure 5 illustrates the performance of models under geometric transformations such as translation and scaling. MSCN consistently exhibits robust performance across all transformations, maintaining stable accuracy despite severe geometric changes. In contrast, other methods experience substantial drops in accuracy, highlighting their vulnerability to such transformations. These results underscore the strength of MSCN in terms of geometrical invariance, demonstrating its potential for real-world applications where robustness to geometric variations is critical.
However, it is important to note that ModelNet40 and ShapeNetPart are derived from CAD models, which are generated in highly idealized environments. In contrast, real-world point clouds are often sparse and only capture surface-level details, which makes them fundamentally different from synthetic datasets. To address this disparity, we extended our experiments to real-world point clouds, enabling a more comprehensive evaluation of our MSCN’s performance in practical scenarios.
V-C Evaluation on Indoor Real-World Datasets: Comparable Performance and Robustness
We conducted experiments on the ScanObjectNN [69] dataset to evaluate the performance of MSCN on real-world indoor point clouds. Table III summarizes the classification results, demonstrating that MSCN achieves performance comparable to the SOTA models in standard evaluation settings. Furthermore, MSCN exhibits remarkable robustness under geometric transformations such as scaling and translation, maintaining stable performance and demonstrating its strength in handling these variations.
However, it is important to note that ScanObjectNN consists of point clouds captured in indoor environments, which are relatively dense and structured. In contrast, point clouds captured in outdoor environments, such as those used in autonomous driving, are significantly sparser and often acquired under harsh conditions, including occlusions and uneven sampling. To address these challenges, we extended our experiments to self-driving vehicle datasets, enabling a more comprehensive evaluation of MSCN’s robustness in real-world scenarios.
TABLE IV: Intra-domain classification results on KITTI [37], nuScenes [38], and PanKyo [42] datasets. The bold black, blue indicate the first and the second best performance on each scenario, respectively.